1. How many grams of magnesium metal will be deposited from a solution that contains Mg2+ions if a current of 0.977 A is applied for 57.8 minutes.

2. How many grams of silver metal will be deposited from a solution that contains Ag+ions if a current of 1.12 A is applied for37.6 minutes.

Answer both please I'm really stuck on these.
Answer #1

From Faraday's first law of electrolysis: The mass of substance deposited is directly proportional to the quantity of electricity passed.

i.e. w q = IT

Where I is the current and t is time

or w = MIt /nF

where M = Molar mass of substance

n = number of electrons involved

F=Faraday's constant = 96485 C /mol

1. Given that, I = 0.977 A = 0.977 C/s

t = 57.8 minutes = 57.8 minutes * 60sec/ 1minute = 3468 s

Molar mass of Mg (M) = 24.305 g/mol

Number of electrons (n) = 2

Now , W = (24.305 g/mol * 0.977 C/s *3468s) / (2* 96485 C/mol) = 0.427 g

2.

Given that, I = 1.12 A = 1.12 C/s

t = 37.6minutes = 37.6 minutes * 60sec/ 1minute = 2256 s

Molar mass of Ag (M) = 107.8682 g/mol

Number of electrons (n) = 1

Now , W = (107.8682 g/mol * 1.12 C/s *2256s) / (1* 96485 C/mol) = 2.82 g